### v1

# A minimal model with stochastically broken reciprocity

Abstract
We introduce a minimal model consisting of a two-body system with stochastically broken reciprocity (i.e., random violation of Newton’s third law) and then investigate its statistical behaviors, including correlation functions, time evolution of probability distribution functions, energy gain, and entropy production. The effective temperature of this two-body system immersed in a thermal bath is derived. Furthermore, we heuristically present an extremely minimal model where the relative motion adheres to the same rules as in classical mechanics, while the effect of stochastically broken reciprocity only manifests in the fluctuating motion of the center of mass.
Keywords: stochastically broken reciprocity, correlation function, probability distribution function, energy gain, entropy production
I Introduction
Newton’s laws of motion are the cornerstone of classical mechanics. Among them, the first law defines inertial frames; the second law establishes the relationship between acceleration and force; and the third law describes the reciprocity of action and reaction forces. Here, reciprocity means that the action force exerted by particle A on particle B and the reaction force exerted by B on A are equal in magnitude, opposite in direction, and collinear with the straight line connecting A and B. This reciprocal nature between action and reaction forces governs not only fundamental microscopic interactions but also the emergent forces between passive particles in equilibrium media [1, 2]. However, reciprocity is found to be broken in non-equilibrium systems where the emergent action and reaction forces are either unequal in magnitude, or out of collinearity [4, 3]. Typically broken phenomena manifest in active systems [11, 5, 6, 7, 8, 9, 10], especially in micro-swimmers [14, 15, 12, 13], active colloids [16, 17, 18, 19, 20], and robotic systems [21, 22, 23, 24].
The strict breaking of reciprocity leads to odd viscosity or elasticity [25, 26, 29, 27, 28, 31, 30], unconventional phase transitions [32, 33, 34, 35, 36, 37, 38], and exotic transport behaviors [39, 40, 41, 42]. Rather than focusing on the aforementioned in-depth discussions on consequences of a strict violation of reciprocity, we aim to explore a more delicate scenario: Newton’s third law holds on average, but reciprocity is broken stochastically. Since the third law holds in average, we expect that the core results of classical mechanics remain valid when expressed in terms of the mean values of relevant physical quantities. Stochastically broken reciprocity, however, plays a role in the fluctuations of these physical quantities. Guided by theoretical interest and pure curiosity, we ask how classical mechanics survives at the level of mean values while stochastic violations imprint themselves on fluctuations.
In this paper, we introduce a minimal model consisting of a two-body system with stochastically broken reciprocity, and then investigate its statistical signatures such as correlation functions, the evolution of probability distribution functions (PDFs), the energy gain and the entropy production, and so on. The remaining content of this paper is organized as follows. In Sec. II, we specifically describe a two-particle system with stochastically broken reciprocity, and derive the dynamic equations of motion for the system’s center of mass (COM) and the two-body relative position based on Newton’s second law. In Sec. III, we computes position-, velocity-, and cross-correlation functions for both modes of motion. In Sec. IV, we derive three Fokker-Planck equations governing the PDFs of the COM motion, the relative motion, and their joint evolution, and use them to track energy and entropy changes. In Sec. V, we immerse the system in a thermal bath and, under overdamped conditions, obtain the Smoluchowski equation for the relative position. In Sec. VI, we present an extremely minimal model in which the relative motion obeys classical mechanics exactly, while the stochastically broken reciprocity merely influence the fluctuating motion of the COM. The final section provides a brief summary and outlook.
II minimal model
Consider a two-body system as depicted in Fig. 1. Particles A and B, with masses $m_{\mathrm{A}}$ and $m_{\mathrm{B}}$, have instantaneous positions $\mathbf{r}_{\mathrm{A}}$ and $\mathbf{r}_{\mathrm{B}}$ measured in an inertial frame. The force exerted by A on B is $\mathbf{F}_{\mathrm{BA}}$; the corresponding reaction force exerted by B on A is $\mathbf{F}_{\mathrm{AB}}$.
Figure 1: Two-body system. Two particles are labeled with A and B, respectively. $\mathbf{F}_{\mathrm{BA}}$ represents the action force exerted by A on B, while $\mathbf{F}_{\mathrm{AB}}$ represents the reaction force exerted by B on A.
We assume that Newton’s third law holds on average. Mathematically, this is expressed as
$\left\langle\mathbf{F}_{\mathrm{AB}}\right\rangle=-\left\langle\mathbf{F}_{\mathrm{BA}}\right\rangle,~{}\mathrm{and}~{}\left\langle\mathbf{F}_{\mathrm{BA}}\right\rangle\parallel\overline{AB}$ (1)
where $\overline{AB}$ denotes the line connecting particles A and B. To render the statement unambiguous, we decompose the instantaneous forces as
$\mathbf{F}_{\mathrm{BA}}=-\nabla\phi(r)+\sqrt{g_{\mathrm{B}}}\bm{\xi}_{\mathrm{B}}(t),$ (2)
$\mathbf{F}_{\mathrm{AB}}=\nabla\phi(r)+\sqrt{g_{\mathrm{A}}}\bm{\xi}_{\mathrm{A}}(t),$ (3)
[…]
$\left\langle\bm{\xi}_{\alpha}(t)\right\rangle=0,$ (4)
and
$\left\langle\bm{\xi}_{\alpha}(t)\bm{\xi}^{\mathrm{T}}_{\beta}(t^{\prime})\right\rangle=\delta_{\alpha\beta}\delta(t-t^{\prime})\mathbf{I},~{}(\alpha,\beta=\mathrm{A,B})$ (5)
with $\delta_{\alpha\beta}$ the Kronecker symbol, $\delta(t-t^{\prime})$ the Dirac delta function, $\mathbf{I}$ the unit tensor. The superscript $T$ represents the transpose operation on vectors or matrices. Obviously, equations (2)–(4) ensure that the average relation (1) is satisfied.
According to Newton’s second law, the equations of motion for the two particles are
$m_{\mathrm{B}}\ddot{\mathbf{r}}_{\mathrm{B}}=-\nabla\phi(r)+\sqrt{g_{\mathrm{B}}}\bm{\xi}_{\mathrm{B}}(t),$ (6)
$m_{\mathrm{A}}\ddot{\mathbf{r}}_{\mathrm{A}}=\nabla\phi(r)+\sqrt{g_{\mathrm{A}}}\bm{\xi}_{\mathrm{A}}(t),$ (7)
Introduce the COM coordinate
$\mathbf{R}\equiv\frac{m_{\mathrm{A}}\mathbf{r}_{\mathrm{A}}+m_{\mathrm{B}}\mathbf{r}_{\mathrm{B}}}{M}$ (8)
with $M\equiv m_{\mathrm{A}}+m_{\mathrm{B}}$ being the total mass of the system. Define the relative position vector between B and A
$\mathbf{r}\equiv\mathbf{r}_{\mathrm{B}}-\mathbf{r}_{\mathrm{A}}.$ (9)
Then Eqs. (6) and (7) yield stochastic differential equations
$\ddot{\mathbf{R}}=\bm{\zeta}_{1}(t),$ (10)
$\ddot{\mathbf{r}}=-\frac{\nabla\phi}{\mu}+\bm{\zeta}_{2}(t),$ (11)
with $\mu\equiv m_{\mathrm{A}}m_{\mathrm{B}}/M$ being the reduced mass of the system. The effective noises are
$\bm{\zeta}_{1}(t)=\frac{\sqrt{g_{\mathrm{A}}}}{M}\bm{\xi}_{\mathrm{A}}(t)+\frac{\sqrt{g_{\mathrm{B}}}}{M}\bm{\xi}_{\mathrm{B}}(t),$ (12)
$\bm{\zeta}_{2}(t)=-\frac{\sqrt{g_{\mathrm{A}}}}{m_{\mathrm{A}}}\bm{\xi}_{\mathrm{A}}(t)+\frac{\sqrt{g_{\mathrm{B}}}}{m_{\mathrm{B}}}\bm{\xi}_{\mathrm{B}}(t).$ (13)
From Eq. (4) we find that these effective noises have vanishing mean:
$\left\langle\bm{\zeta}_{i}(t)\right\rangle=0,~{}(i=1,2).$ (14)
In addition, from Eq. (5) we obtain their correlations
$\left\langle\bm{\zeta}_{i}(t)\bm{\zeta}^{\mathrm{T}}_{j}(t^{\prime})\right\rangle=g_{ij}\delta(t-t^{\prime})\mathbf{I},~{}(i,j=1,2)$ (15)
where the coefficients are explicitly expressed as
$g_{11}=\frac{g_{\mathrm{A}}+g_{\mathrm{B}}}{M^{2}},$ (16)
$g_{12}=g_{21}=\frac{1}{M}(\frac{g_{\mathrm{B}}}{m_{\mathrm{B}}}-\frac{g_{\mathrm{A}}}{m_{\mathrm{A}}}),$ (17)
$g_{22}=\frac{g_{\mathrm{A}}}{m_{\mathrm{A}}^{2}}+\frac{g_{\mathrm{B}}}{m_{\mathrm{B}}^{2}}.$ (18)
Stochastic differential equations (10) and (11) combining with (14) and (15) govern the COM motion and the relative motion between B and A.
III correlation functions
The observable consequences of stochastically broken reciprocity are encoded in the correlation structure of the system. In what follows we examine the position-position, velocity-velocity, and position-velocity correlation functions for both the COM motion and the relative motion of the two particles. To be concrete, we restrict the discussion to one spatial dimension and take the interaction potential to be of harmonic form.
III.1 Correlation functions for the COM motion
In one-dimensional situation, let $X$ and $V$ denote the position and velocity of the COM. From Eq. (10) together with Eqs. (14) and (15), we obtain the equations governing the COM motion
$\dot{X}=V,$ (19)
$\dot{V}=\zeta_{1}(t),$ (20)
where the Gaussian white noise ${\zeta}_{1}(t)$ satisfies
$\left\langle{\zeta}_{1}(t)\right\rangle=0,$ (21)
and
$\left\langle{\zeta}_{1}(t){\zeta}_{1}(t^{\prime})\right\rangle=g_{11}\delta(t-t^{\prime}).$ (22)
By integrating Eq. (20), we immediately obtain the COM velocity
$V=V_{0}+\int_{0}^{t}\zeta_{1}(\tau)\mathrm{d}\tau,$ (23)
where $V_{0}$ represents the initial velocity of the COM. By integrating the above equation again and exchanging the order of integration for double integral, we can obtain the COM displacement
$X-X_{0}=V_{0}t+\int_{0}^{t}(t-\tau)\zeta_{1}(\tau)\mathrm{d}\tau,$ (24)
with $X_{0}$ the initial position of the COM. Considering Eq. (21), we immediately obtain $\left\langle V\right\rangle=\left\langle V_{0}\right\rangle$ and $\left\langle X-X_{0}\right\rangle=\left\langle V_{0}\right\rangle t$ from Eqs. (23) and (24). Hence, on average, the centre of mass executes uniform rectilinear motion.
We now calculate velocity autocorrelation function. Squaring Eq. (23) and considering Eq. (21), we have
$\langle V^{2}\rangle=\langle V_{0}^{2}\rangle+\left\langle\left[\int_{0}^{t}\zeta_{1}(\tau)\mathrm{d}\tau\right]^{2}\right\rangle.$ (25)
[…]
$\langle V^{2}\rangle=\langle V_{0}^{2}\rangle+g_{11}t.$ (26)
Repeating the similar procedure, we can achieve the velocity-displacement correlation
$\left\langle(X-X_{0})V\right\rangle=\left\langle V^{2}_{0}\right\rangle t+\frac{g_{11}}{2}t^{2}$ (27)
and the displacement (related to position) autocorrelation function
$\left\langle(X-X_{0})^{2}\right\rangle=\left\langle V^{2}_{0}\right\rangle t^{2}+\frac{g_{11}}{3}t^{3}.$ (28)
The linear, quadratic, and cubic time dependences in correlation functions (26)–(28) reflect the effect of stochastically broken reciprocity. When $g_{11}=0$, these results reduce to the familiar situation of uniform rectilinear motion in classical mechanics.
III.2 Correlation functions for the relative motion of B with respect to A
In one-dimensional situation, the relative position and velocity of B with respect to A are denoted as $x$ and $v$, respectively. According to Eqs. (11), (14) and (15), the equation of relative motion of B with respect to A can be expressed as
$\dot{x}=v,$ (29)
$\dot{v}=-\phi^{\prime}/\mu+\zeta_{2}(t),$ (30)
where $\phi^{\prime}\equiv\mathrm{d}\phi/\mathrm{d}x$, and $\zeta_{2}(t)$ is a Gaussian white noise satisfying
$\left\langle{\zeta}_{2}(t)\right\rangle=0,$ (31)
and
$\left\langle{\zeta}_{2}(t){\zeta}_{2}(t^{\prime})\right\rangle=g_{22}\delta(t-t^{\prime}).$ (32)
For simplicity, we take the harmonic potential $\phi=\frac{1}{2}\mu\omega^{2}x^{2}$ so that Eq. (30) reduces to
$\dot{v}=-\omega^{2}x+\zeta_{2}(t).$ (33)
Equations (29) and (33) describe a linear, damp-free harmonic oscillator driven by the stochastic force $\zeta_{2}(t)$. Following the standard procedure we achieve the solution
$\left(\begin{array}[]{c}v\\ \omega x\\ \end{array}\right)=U\left(\begin{array}[]{c}v_{0}\\ \omega x_{0}\\ \end{array}\right)+\int_{0}^{t}\zeta_{2}(\tau)\left(\begin{array}[]{c}\cos\omega(t-\tau)\\ \sin\omega(t-\tau)\\ \end{array}\right)\mathrm{d}\tau$ (34)
with the rotation matrix $U=\left(\begin{array}[]{cc}\cos\omega t&-\sin\omega t\\ \sin\omega t&\cos\omega t\\ \end{array}\right)$. Here $v_{0}$ and $x_{0}$ denote respectively the initial relative velocity and position of B with respect to A. Taking the average on Eq. (34) and using Eq. (31), we immediately obtain
$\left(\begin{array}[]{c}\langle v\rangle\\ \omega\langle x\rangle\\ \end{array}\right)=U\left(\begin{array}[]{c}\langle v_{0}\rangle\\ \omega\langle x_{0}\rangle\\ \end{array}\right),$ (35)
so the averaged motion reduces to the familiar undamped harmonic oscillation in classical mechanics.
We now evaluate the correlation functions. Define a covariance matrix
${\sigma}(t)\equiv\left(\begin{array}[]{cc}\langle v^{2}(t)\rangle&\omega\langle v(t)x(t)\rangle\\ \omega\langle x(t)v(t)\rangle&\omega\langle x^{2}(t)\rangle\\ \end{array}\right).$ (36)
Using Eqs. (31),(32), and the solution (34), one obtains
$\sigma(t)=U\sigma(0)U^{\mathrm{T}}+\frac{g_{22}}{2\omega}\left(\begin{array}[]{cc}\omega t+\frac{\sin 2\omega t}{2}&\frac{1-\cos 2\omega t}{2}\\ \frac{1-\cos 2\omega t}{2}&\omega t-\frac{\sin 2\omega t}{2}\\ \end{array}\right)$ (37)
The second term on the right-handed side of the above equation is the signature of stochastically broken reciprocity. When $g_{22}=0$, Eq. (37) returns to the familiar oscillator correlations in classical mechanics.
IV PDFs, energy and entropy
In this section, we will derive three Fokker-Planck equations characterizing the evolutions of PDFs for the COM motion, the two-body relative motion, and their joint evolution, respectively. With these PDFs we then examine how stochastically broken reciprocity gives rise to both energy gain and entropy production in the system.
IV.1 PDF for the COM motion
Since equations of motion of the COM, (19) and (20), are stochastic, we may define a PDF $\rho(X,V,t)$ such that $\rho(X,V,t)\mathrm{d}X\mathrm{d}V$ represents the probability of finding the COM position and velocity at time $t$ within the small rectangle with edges $\mathrm{d}X$ and $\mathrm{d}V$ centered at the phase point $(X,V)^{\mathrm{T}}$.
Deriving the evolution of the PDF from the stochastic equations of motion is a standard topic in modern statistical physics textbooks [43]. Following the method outlined in Ref. [43], we can straightforwardly derive the Fokker-Planck equation
$\frac{\partial\rho}{\partial t}=-V\frac{\partial\rho}{\partial X}+\frac{g_{11}}{2}\frac{\partial^{2}\rho}{\partial V^{2}}$ (38)
which describes the evolution of the PDF of the COM.
IV.2 PDF for the two-body relative motion
The two-body relative motion is governed by equations (29) and (30). we may define a PDF $\varrho(x,v,t)$ such that $\varrho(x,v,t)\mathrm{d}x\mathrm{d}v$ represents the probability of finding the relative position and velocity at time $t$ within the small rectangle with edges $\mathrm{d}x$ and $\mathrm{d}v$ centered at the phase point $(x,v)^{\mathrm{T}}$. Following the method in Ref. [43], we can easily derive the Fokker-Planck equation
$\frac{\partial\varrho}{\partial t}=-v\frac{\partial\varrho}{\partial x}+\frac{\phi^{\prime}}{\mu}\frac{\partial\varrho}{\partial v}+\frac{g_{22}}{2}\frac{\partial^{2}\varrho}{\partial v^{2}},$ (39)
which describes the evolution of the PDF corresponding to the two-body relative motion.
IV.3 Joint PDF for the COM motion and the relative motion
We define a joint PDF $f(X,V,x,v,t)$ such that $f(X,V,x,v,t)\mathrm{d}\Gamma$ represents the probability of finding the COM position and velocity, and the relative position and velocity of B with respect to A, within the small element $\mathrm{d}^{4}\Gamma\equiv\mathrm{d}X\mathrm{d}V\mathrm{d}x\mathrm{d}v$ centered at the phase point $\bm{\Gamma}\equiv(X,V,x,v)^{\mathrm{T}}$. Since Eq. (17) implies $\langle\zeta_{1}(t)\zeta_{2}(t)\rangle\neq 0$, the joint PDF $f(X,V,x,v,t)$ usually differs from $\rho(X,V,t)\varrho(x,v,t)$. We may derive the evolution of the joint PDF following the standard stochastic method in Ref. [44].
Considering the transformation relations (12) and (13), we can rewrite Eqs. (19), (20), (29) and (30) in a compact form as
$\mathrm{d}\bm{\Gamma}=\bm{\Lambda}\mathrm{d}t+\Omega\mathrm{d}\mathbf{W}(t)$ (40)
where the deterministic phase velocity $\bm{\Lambda}$ and the noise strength matrix $\Omega$ can be explicitly expressed as
$\bm{\Lambda}=(V,0,v,-\phi^{\prime}/\mu)^{\mathrm{T}}$ (41)
and
$\Omega=\left(\begin{array}[]{cc}0&0\\ \sqrt{g_{\mathrm{A}}}/M&\sqrt{g_{\mathrm{B}}}/M\\ 0&0\\ -\sqrt{g_{\mathrm{A}}}/m_{\mathrm{A}}&\sqrt{g_{\mathrm{B}}}/m_{\mathrm{B}}\\ \end{array}\right).$ (42)
The noise term $\mathrm{d}\mathbf{W}(t)=\left(\xi_{\mathrm{A}}(t)\mathrm{d}t,\xi_{\mathrm{B}}(t)\mathrm{d}t\right)^{\mathrm{T}}$ represents two-variable Wiener process.
According to the method in Ref. [44], we can write the corresponding Fokker-Planck equation as
$\frac{\partial f}{\partial t}=-\nabla_{\Gamma}\cdot\mathbf{J},$ (43)
with the flux
$\mathbf{J}$ $=$ $\bm{\Lambda}f-\frac{1}{2}\Omega\Omega^{\mathrm{T}}\nabla_{\Gamma}f$ (48)
$=$ $\left(\begin{array}[]{c}Vf\\ -\frac{g_{11}}{2}\frac{\partial f}{\partial V}-\frac{g_{12}}{2}\frac{\partial f}{\partial v}\\ vf\\ -\frac{\phi\prime}{\mu}f-\frac{g_{12}}{2}\frac{\partial f}{\partial V}-\frac{g_{22}}{2}\frac{\partial f}{\partial v}\\ \end{array}\right)$
where $g_{11}$, $g_{12}$ and $g_{33}$ are coefficients shown in Eqs. (16)–(18). Note that the symbol $\nabla_{\Gamma}$ represents the gradient operator on the phase space $\{\bm{\Gamma}\equiv(X,V,x,v)^{\mathrm{T}}\}$.
With the above evolution equations of the joint PDF, we can easily verify Eqs. (38) and (39) using $\rho(X,V,t)=\int f(X,V,x,v,t)\mathrm{d}x\mathrm{d}v$, $\varrho(x,v,t)=\int f(X,V,x,v,t)\mathrm{d}X\mathrm{d}V$, and the Stokes’ theorem familiar in multivariable calculus.
IV.4 Energy gain
Based on the experience of stochastic thermodynamics [45, 46, 47], we may define the energy as the average mechanical energy in classical mechanics, which reads
$E$ $=$ $\left\langle\frac{M}{2}V^{2}+\frac{\mu}{2}v^{2}+\phi(x)\right\rangle$ (49)
$=$ $\int\left[\frac{M}{2}V^{2}+\frac{\mu}{2}v^{2}+\phi(x)\right]f\mathrm{d}^{4}\Gamma$
$=$ $\frac{M}{2}\int V^{2}\rho\mathrm{d}X\mathrm{d}V+\int\left[\frac{\mu}{2}v^{2}+\phi(x)\right]\varrho\mathrm{d}x\mathrm{d}v$
Using Eq. (38) and the Stokes’ theorem, and assuming vanishing boundary integrals at infinity, we can obtain
$\int V^{2}\frac{\partial\rho}{\partial t}\mathrm{d}X\mathrm{d}V=g_{11}.$ (50)
Similarly, using Eq. (39) and the Stokes’ theorem, and assuming vanishing boundary integrals at infinity, we can obtain
$\int\left[\frac{\mu}{2}v^{2}+\phi(x)\right]\frac{\partial\varrho}{\partial t}\mathrm{d}x\mathrm{d}v=\frac{\mu g_{22}}{2}.$ (51)
Therefore, we eventually arrive at
$\frac{\mathrm{d}E}{\mathrm{d}t}$ $=$ $\frac{M}{2}\int V^{2}\frac{\partial\rho}{\partial t}\mathrm{d}X\mathrm{d}V+\int\left[\frac{\mu}{2}v^{2}+\phi(x)\right]\frac{\partial\varrho}{\partial t}\mathrm{d}x\mathrm{d}v$ (52)
$=$ $\frac{Mg_{11}+\mu g_{11}}{2}\geq 0,$
which implies that the system is experiencing an energy gain.
IV.5 Entropy production
Following the concept of stochastic thermodynamics [46], we define the trajectory entropy as
$s=-\ln f,$ (53)
where the prefactor $k_{B}$ has been set to unity. The entropy is then regarded as the average of the trajectory entropy, which reads
$S=\langle s\rangle=-\int f\ln f\mathrm{d}^{4}\Gamma.$ (54)
The rate of entropy change is
$\frac{\mathrm{d}S}{\mathrm{d}t}=-\int\frac{\partial f}{\partial t}(\ln f+1)\mathrm{d}^{4}\Gamma.$ (55)
Substituting Eq. (43) into the above equation, then using the Stokes’ theorem, and assuming vanishing boundary integrals at infinity, we obtain
$\frac{\mathrm{d}S}{\mathrm{d}t}=\int\mathbf{J}\cdot\nabla_{\Gamma}(\ln f+1)\mathrm{d}^{4}\Gamma.$ (56)
Substituting the expression (48) for the flux $\mathbf{J}$ into the above equation, we can derive
$\frac{\mathrm{d}S}{\mathrm{d}t}=\frac{1}{2}\int\left(\sum_{ij}g_{ij}s_{i}s_{j}\right)f\mathrm{d}^{4}\Gamma.$ (57)
with $s_{1}\equiv{\partial s}/{\partial V}$ and $s_{2}\equiv{\partial s}/{\partial v}$. From Eqs. (16)–(18), we calculate $g_{11}g_{22}-g_{12}^{2}={g_{\mathrm{A}}g_{\mathrm{B}}}/{m_{\mathrm{A}}^{2}m_{\mathrm{B}}^{2}}\geq 0$. Thus, the integrand in Eq. (57) is a positive-definite form, and hence ${\mathrm{d}S}/{\mathrm{d}t}\geq 0$. The stochastically broken reciprocity results in the entropy production of the two-body system.
V Two-body system immersed in a thermal bath
We further ask what will happen if we place the two-body system mentioned above in a thermal bath at a constant temperature $T$. For simplicity, we set $k_{B}$ to unity and discuss the overdamped situation.
The equations of motion can be expressed as
$\gamma_{\mathrm{A}}\dot{\mathbf{r}}_{\mathrm{A}}=\nabla\phi(r)+\sqrt{g_{\mathrm{A}}}\bm{\xi}_{\mathrm{A}}(t)+\sqrt{2\gamma_{\mathrm{A}}T}\bm{\xi}_{\mathrm{TA}}(t),$ (58)
$\gamma_{\mathrm{B}}\dot{\mathbf{r}}_{\mathrm{B}}=-\nabla\phi(r)+\sqrt{g_{\mathrm{B}}}\bm{\xi}_{\mathrm{B}}(t)+\sqrt{2\gamma_{\mathrm{B}}T}\bm{\xi}_{\mathrm{TB}}(t),$ (59)
where $\gamma_{\mathrm{A}}$ and $\gamma_{\mathrm{B}}$ are the damping coefficients of particles A and B, respectively. The terms $\sqrt{2\gamma_{\mathrm{A}}T}\bm{\xi}_{\mathrm{TA}}(t)$ and $\sqrt{2\gamma_{\mathrm{B}}T}\bm{\xi}_{\mathrm{TB}}(t)$ represent thermal noises due to the bath, which are assumed to be Gaussian white noise. Additionally, we assume that $\bm{\xi}_{\mathrm{A}}(t)$, $\bm{\xi}_{\mathrm{B}}(t)$, $\bm{\xi}_{\mathrm{TA}}(t)$, and $\bm{\xi}_{\mathrm{TB}}(t)$ are independent of each other.
From Eq. (58) and (59), we can derive the equation of two-body relative motion
$\dot{\mathbf{r}}=-\frac{\nabla\phi}{\nu}+\bm{\chi}(t),$ (60)
where the reduced damping coefficient $\nu\equiv{\gamma_{\mathrm{A}}\gamma_{\mathrm{B}}}/({\gamma_{\mathrm{A}}+\gamma_{\mathrm{B}}})$, and the noise term is given by $\bm{\chi}(t)=({\sqrt{g_{\mathrm{B}}}}/{\gamma_{\mathrm{B}}})\bm{\xi}_{\mathrm{B}}(t)+\sqrt{{2T}/{\gamma_{\mathrm{B}}}}\bm{\xi}_{\mathrm{TB}}(t)-({\sqrt{g_{\mathrm{A}}}}/{\gamma_{\mathrm{A}}})\bm{\xi}_{\mathrm{A}}(t)-\sqrt{{2T}/{\gamma_{\mathrm{A}}}}\bm{\xi}_{\mathrm{TA}}(t)$. It is not hard to verify that
$\left\langle\bm{\chi}(t)\right\rangle=0,$ (61)
and
$\left\langle\bm{\chi}(t)\bm{\chi}^{\mathrm{T}}(t^{\prime})\right\rangle=2D\mathbf{I}\delta(t-t^{\prime})$ (62)
with $D\equiv{T}/\nu+{{g_{\mathrm{A}}}}/{2\gamma_{\mathrm{A}}^{2}}+{{g_{\mathrm{B}}}}/{2\gamma_{\mathrm{B}}^{2}}$. According to the Einstein relation [43], we may define an effective temperature
$T_{e}\equiv\nu D={T}+\frac{\nu}{2}\left(\frac{{g_{\mathrm{A}}}}{\gamma_{\mathrm{A}}^{2}}+\frac{{g_{\mathrm{B}}}}{\gamma_{\mathrm{B}}^{2}}\right),$ (63)
which is clearly larger than the bath temperature $T$.
Following the method in Ref. [43], we can easily derive the Smoluchowski equation corresponding to Eqs. (60)–(62). The PDF $P(\mathbf{r},t)$ for the two-body relative motion satisfies
$\frac{\partial P}{\partial t}=D\nabla\cdot\left[\frac{(\nabla\phi){P}}{T_{e}}+\nabla P\right].$ (64)
From the above equation, we observe that a steady state exists. The steady-state PDF has the Boltzmann form as
$P\propto\exp\left\{-\frac{\phi(x)}{T_{e}}\right\}.$ (65)
VI Extremely minimal model
In the minimal model mention above, $\bm{\xi}_{\mathrm{A}}(t)$ and $\bm{\xi}_{\mathrm{B}}(t)$ are assumed to be independent of each other. We may heuristically consider an extreme situation in which $\bm{\xi}_{\mathrm{A}}(t)$ and $\bm{\xi}_{\mathrm{B}}(t)$ are not independent such that $\bm{\zeta}_{2}(t)=0$. In this sense, the two-body system is referred to as the extremely minimal model.
By setting Eq. (13) to be vanishing, we obtain a necessary condition: $\frac{\sqrt{g_{\mathrm{A}}}}{m_{\mathrm{A}}}\bm{\xi}_{\mathrm{A}}(t)=\frac{\sqrt{g_{\mathrm{B}}}}{m_{\mathrm{B}}}\bm{\xi}_{\mathrm{B}}(t)$, which enlightens us to assume $\sqrt{g_{\mathrm{A}}}\bm{\xi}_{\mathrm{A}}(t)=m_{\mathrm{A}}\sqrt{g}\bm{\xi}(t)$ and $\sqrt{g_{\mathrm{B}}}\bm{\xi}_{\mathrm{B}}(t)=m_{\mathrm{B}}\sqrt{g}\bm{\xi}(t)$ where $g$ is a positive constant quantity. To guarantee this assumption, we need to revisit the main equations in Sec. II. Equations (2) and (3) are rewritten as
$\mathbf{F}_{\mathrm{BA}}=-\nabla\phi(r)+m_{\mathrm{B}}\sqrt{g}\bm{\xi}(t),$ (66)
$\mathbf{F}_{\mathrm{AB}}=\nabla\phi(r)+m_{\mathrm{A}}\sqrt{g}\bm{\xi}(t).$ (67)
Equations (4) and (5) are replaced with
$\left\langle\bm{\xi}(t)\right\rangle=0,$ (68)
and
$\left\langle\bm{\xi}(t)\bm{\xi}^{\mathrm{T}}(t^{\prime})\right\rangle=\delta(t-t^{\prime})\mathbf{I}.$ (69)
The equations governing the COM motion and the relative motion of B with respect to A are revised to
$\ddot{\mathbf{R}}=\sqrt{g}\bm{\xi}(t),$ (70)
$\ddot{\mathbf{r}}=-\frac{\nabla\phi}{\mu}.$ (71)
Thus, this system is more concise, as the relative motion follows the same rule as in classical mechanics. The COM maintains uniform rectilinear motion on average. The stochastically broken reciprocity only affects the fluctuating motion of the COM. The main conclusions in Secs.IIIA and IVA remain unchanged. The other conclusions need to be reexamined in detail.
VII conclusion
In the above discussion, we have proposed a minimal model consisting of a two-body system with stochastically broken reciprocity. Guided by Newton’s second law, we have derived two stochastic differential equations that govern the COM motion and the two-body relative motion. Based on these equations of motion, we have obtained the position-, velocity-, and cross-correlation functions for the COM motion and the two-body relative motion, respectively. Additionally, we have derived three Fokker-Planck equations, which respectively characterize the evolution of PDFs for the COM motion, the two-body relative motion, and the joint evolution of these two motions. Using these Fokker-Planck equations, we have analyzed the energy gain and entropy production behaviors in this two-body system arising from the stochastically broken reciprocity. We have further explored the two-body system in a constant-temperature thermal bath and determined the effective temperature of the system under the overdamped condition. The corresponding Smoluchowski equation, which describes the evolution of the PDF of the two-body relative position, yields the Boltzmann distribution with the effective temperature in the steady state. Finally, we have introduced an extremely minimal model where the relative motion adheres to the laws of classical mechanics, and the effect of stochastically broken reciprocity is solely manifested in the fluctuating motion of the COM.
Before concluding this paper, we discuss three prospective issues for future research:
1) Experimental verification. A key question is whether the minimal model with stochastically broken reciprocity can be experimentally tested. As we have predicted, the correlation functions (Eqs. 26, 27, and 28) contain linear, quadratic, and cubic time-scaling terms, respectively. These scaling laws could be examined by observing the fluctuating motion of the COM in experiments. However, we regret that we cannot currently specify a real-world system where these effects might be observed, leaving this as a challenge for experimental physicists.
2) Extension to many-body systems. The framework of the present minimal model can be extended to many-body systems, where pairwise interactions are assumed to stochastically violate Newton’s third law following the same mechanism as in Eqs. 2 and 3. If no additional constraints are imposed on the entire system, the COM motion is expected to follow a rule similar to Eq. 10, meaning the correlation behaviors for the COM motion (discussed in Sec. IIIA) would remain valid. However, the motion of particles relative to the COM would be far more complex than in the two-body case, warranting further investigation.
3) Implication for fundamental forces. We have believed that the fundamental forces in classical mechanics strictly obey Newton’s third law. Experimental observations by scientists typically focus on the relationships between the mean values of physical quantities. From a theoretical perspective, however, we cannot entirely exclude the possibility that certain fundamental forces may, on average, satisfy Newton’s third law but stochastically violate it. Introducing such stochastic violation into fundamental interactions could have profound implications for fields ranging from quantum mechanics and particle physics to cosmology, though this remains a speculative direction requiring rigorous theoretical and experimental scrutiny.
ACKNOWLEDGMENTS
The author is grateful to the support from the National Natural Science Foundation of China (Grant No. 12475032).